OpenStudy (anonymous):

Find the length of the diameter of a sphere with a volume of 1767.15 ft3.

OpenStudy (anonymous):

so far I have gotten: 1767.15=4/3πr^3 1325.36=πr^3 421.88=r^3 but i dont know how to do the last part.

OpenStudy (tkhunny):

Have you considered the cube root? \(r = 421.88^{1/3}\)

OpenStudy (anonymous):

it has to be a number without a cubed root.. the options are: 15 ft 20.5 ft 41 ft 7.5 ft

OpenStudy (tkhunny):

Oh, well, in that case, have you considered the cube root? \(r=421.88^{1/3}\)

OpenStudy (anonymous):

i dont know how to do cube root.. when using a calculator do you just put in 421.88 to the 1/3?

OpenStudy (tkhunny):

Yes. There should a button like \(x^{y}\). You can also use logarithms.

OpenStudy (anonymous):

so the answer would be d right?

OpenStudy (tkhunny):

I get 7.5000

OpenStudy (anonymous):

okay.. thank you :)

OpenStudy (whpalmer4):

If you have a list of possible answers, and can't figure out how to do a cube root, you could always try cubing the various answers and see which one gives you the number for which you couldn't find the cube root :-)

OpenStudy (whpalmer4):

In fact, if your answers are far enough apart, you can probably just estimate it. The cube root of 421.88 is the number x such that x*x*x = 421.88. Well, let's try 10, that's easy to cube: 10*10*10 = 1000. Oops, too big. Well, 3 of the answers are larger than 10, and we know 10 is too big, so that just leaves the 7.5. As a quick check, 7*7*7 = 49 * 7 which is about 50 * 7 = 350. That's pretty close to 421, and 7.5*7.5*7.5 would obviously be a bit larger.